# A New Model-Independent Precision Test of General Relativity using LISA Bright Standard Sirens

### v1

@@ Abstract
The upcoming Laser Interferometer Space Antenna (LISA), set for launch in the mid-2030s, will enhance our capability to probe the universe through gravitational waves (GWs) emitted from binary black holes (BBHs) across a broad range of cosmological distances. LISA is projected to observe three classes of BBHs: massive BBHs (MBBHs), extreme mass-ratio inspirals (EMRIs), and stellar mass BBHs. This study focuses on MBBHs, which are anticipated to occur in gas-rich environments conducive to producing powerful electromagnetic (EM) counterparts. This capability to precisely localize MBH events in the sky positions them as excellent candidates for bright sirens. By combining GW luminosity distance measurements from these bright sirens with Baryon Acoustic Oscillation (BAO) measurements derived from galaxy clustering and sound horizon measurements from the Cosmic Microwave Background (CMB), we propose a data-driven model-independent method to reconstruct deviations in the variation of the effective Planck mass (in conjunction with the Hubble constant) as a function of cosmic redshift. Using this multi-messenger technique, we achieve precise measurements of deviations in the effective Planck mass variation with redshift (z), with a precision ranging from approximately $2.4\%$ to $7.2\%$ from redshift $z=1$ to $z=6$ with a single event. Additionally, we achieved a measurement of the Hubble constant with a precision of about $1.3\%$, accounting for variations in the effective Planck mass over 4 years of observation time ($T_{\mathrm{obs}}$). This assumes that EM counterparts are detected for $75\%$ of the events. This precision improves with observation time as $T_{\mathrm{obs}}^{-1/2}$. This approach not only has the potential to reveal deviations from General Relativity but also to significantly expand our understanding of the universe’s fundamental physical properties.
@@ 1 Introduction
However, it is the forthcoming space-based observatory, the Laser Interferometer Space Antenna (LISA) [17, 18, 19], that is set to revolutionize our understanding by observing GWs from binary black holes (BBHs) over unprecedented cosmological distances. LISA is designed to explore the milliHertz frequency range, which is rich with sources such as massive BBHs (MBBHs) [20], extreme mass ratio inspirals (EMRIs) [21, 22], and stellar mass BHBs [23] each with unique observational characteristics and expected in different redshift ranges. This ability to probe such diverse astrophysical phenomena is crucial because ground-based detectors are hindered by unavoidable noise artifact, which prevents them from accessing this lower frequency domain. Scheduled for launch in the mid-2030s, LISA will detect these sources up to very high redshifts with unparalleled precision [17]. Its ability to observe with unique observational characteristics across different redshift ranges positions it as a pivotal cosmological tool [24]. This extensive range allows for the mapping of the universe’s expansion from local to very distant realms. Notably, MBBHs, which are anticipated to form in gas-rich environments, may generate significant electromagnetic (EM) counterparts through various processes such as jets, disk winds, or accretion [25, 26, 27, 28, 29, 30]. The detection of these EM emissions not only enables precise localization of MBH events in the sky but also makes them exceptionally valuable as bright sirens for gravitational studies. The combination of GWs and EM observations could lead to breakthroughs in multi-messenger astronomy, further enhancing our understanding of the universe’s structure and evolution.
@@
Along with these, the investigation of modified gravity theories can gain insights through the propagation speed and luminosity distance of GWs and EM signals. Various factors, including the mass of gravitons, a frictional term related to the changing Planck mass, and anisotropic source terms, influence deviations from expected results [31, 32, 33, 34, 35, 36, 37]. Accurate measurements of these deviations provide a framework for rigorously testing alternate gravitational theories. Utilizing bright sirens is a particularly effective method for evaluating parameters like the GWs propagation speed, the mass of gravitons, and frictional effects denoted by $\gamma(z)$ (see in Equation.2.2). The observation of an EM counterpart roughly 1.7 seconds following the binary neutron star merger GW170817 enabled the establishment of stringent limits on the speed of GWs and the graviton’s mass [6, 38, 39]. Some theories suggest variations in GWs speed, especially within frequency bands monitored by LISA, contrasting with those covered by LIGO [40]. For the purpose of this study, we assume that the speed of GWs is equivalent to that of light.
@@
The literature offers various model-dependent parameterizations of the friction term $\gamma(z)$, including the popular ($\Xi_{0},n$) model [35], which has been explored for its detectability with observatories like LVK [41, 42] and LISA [43]. This particular parameterization utilizes two positive parameters, $\Xi_{0}$ and $n$, where a $\Xi_{0}$ value of 1 aligns with GR. Various models of modified gravity propose different behaviors for this frictional term under diverse conditions. Scalar-Tensor theories, for example, introduce a scalar field that significantly influences cosmological dynamics. Notable within this category are Brans–Dicke theory [44], f(R) gravity [45], and covariant Galileon models [46], all part of the broader Horndeski framework. Extending beyond Horndeski, Degenerate Higher Order Scalar-Tensor (DHOST) theories [47, 48, 49] represent a comprehensive category of scalar-tensor models that propagate a solitary scalar degree of freedom along with the massless graviton’s helicity-2 mode. Other notable theories include f(Q) gravity, f(T) gravity [50], minimal bigravity [51], and models incorporating extra dimensions [52, 53, 54].
@@
These theories demonstrate varying deviations from GR that are not fully captured by the power-law $\Xi_{0}$ and $n$ parameterization, particularly at the extremes of the redshift spectrum, where the fitting formula for $\gamma(z)$ may lose accuracy [31]. Moreover, the integrated effect of $\gamma(z)$ in the real universe may deviate from a simple power-law, underscoring the necessity for model-independent methods to test GR and measure $\gamma(z)$ based on observational data. To address these issues, we propose a novel model-independent approach through a function designated as $\mathcal{F}(z)$, defined in Equation 2.4. This function is crafted to quantify any departures from GR by comparing GWs luminosity distances $\mathrm{D_{l}^{GW}(z)}$ with their EM counterparts $\mathrm{D_{l}^{EM}(z)}$. In the context of standard GR, $\mathcal{F}(z)$ should theoretically equal 1, providing a clear metric to gauge discrepancies attributable to alternative theories of gravity.
@@
Figure 1: This diagram outlines the methodology for measuring three critical length scales: the sound horizon ($r_{s}$), the BAO scale ($\theta_{BAO}$), and the GWs luminosity distance ($D_{l}^{GW}$). These scales are derived from distinct observational methods, including CMB measurements, galaxy correlation studies, and detections of GWs. The interplay between the sound horizon and the BAO scale establishes the EM angular diameter distance ($D^{\text{EM}}_{a}$). Utilizing the distance duality relation, these measurements facilitate a redshift-dependent empirical verification of GR, as expressed in Equation 2.9. This schematic provides a visual framework for understanding the multi-messenger approach in extracting these scales using LISA’s bright sirens.
@@ 2 Non-parametric reconstruction of deviation from General Relativity: Formalism
where $h_{(+,\times)}$ represents the strain of the GWs in the plus ($+$) and cross ($\times$) polarizations, the prime denotes a derivative with respect to conformal time ($\eta$), and $\mathcal{H}$ is the Hubble parameter in comoving coordinates. This equation forms the basis for analyzing GW propagation and testing the validity of GR. Alterations to this equation in the context of modified gravity theories are expressed as
@@
where $\gamma(z)$ represents a frictional term reflecting the influence of modified dynamics, $c_{GW}$ denotes the speed of GWs propagation, $m_{GW}$ is the graviton mass, $a$ is the scale factor, and $\Pi_{(+,\times)}$ is the anisotropic stress term.
Under standard GR assumptions, all the additional parameters such as $\gamma(z)$, $m_{GW}$, and $\Pi_{(+,\times)}$ are set to zero, except for $c_{GW}$, which is equal to the speed of light (denoted by $c$). When testing the deviations from GR, these parameters are crucial to ascertain any potential discrepancies. Observations from the GW170817 event have placed strict limits, confirming that the graviton mass ($m_{GW}$) is effectively zero, and that the speed of GWs propagation ($c_{GW}$) aligns with the speed of light [38, 39]. Although some theories suggest variations in the speed of GWs outside the LVK observational frequency range [40], for the purpose of this analysis, we maintain the assumption that the speed of GWs is equivalent to that of light. Furthermore, changes in the anisotropic stress term ($\Pi_{+,\times}$) can affect the phase of binary waveforms. Recent observations, especially of GW events like GW150914 and GW151226, have begun to constrain these potential deviations, although the constraints remain relatively broad [55]. The friction term, denoted as $\gamma(z)$, significantly affects the amplitude of GWs from cosmological sources. This alteration suggests that the luminosity distance to GWs, measured via standard sirens, could differ from distances determined through EM probes like standard candles, CMB, or BAO. This difference potentially serves as a distinct indicator of modified gravity theories. Hence, the GW luminosity distance, $\mathrm{D_{l}^{GW}}$, is linked to the EM luminosity distance, $\mathrm{D_{l}^{EM}}$, through an exponential relationship with the exponent being the integral of $\mathrm{\gamma(z^{\prime})/(1+z^{\prime})}$ [34]
@@
here, $\mathcal{F}(z)$ encapsulates any deviations from GR as a function of redshift, aiding in the identification of modified gravity effects through observed GWs and EM luminosity distances introduced in the paper[56] . Alternatively, a parametric approach that employs $\Xi_{0}$ and $n$ is used to describe a power-law modification in GW propagation over redshift[57, 58, 43]
@@
GWs offer a unique method to measure luminosity distances to binary systems [59]. While standard $\Lambda$CDM cosmology can also determine these distances, it relies on model-dependent assumptions. In contrast, a model-independent, data-driven method involves using EM luminosity distances derived from BAO’s angular diameter distance, $D^{\rm EM}_{A}(z)$, calculated at various redshifts
$r_{s}=\int_{z_{d}}^{\infty}\frac{dz\,c_{s}(z)}{H_{0}\sqrt{\Omega_{m}(1+z)^{3}+\Omega_{\Lambda}+\Omega_{r}(1+z)^{4}}},$ (2.6)
where $\Omega_{m}$ is the matter density parameter, $\Omega_{\Lambda}$ is the dark energy density parameter, $\Omega_{r}$ is the radiation density parameter, and $H_{0}$ is the Hubble constant, which denotes the current rate of expansion of the universe. By integrating this data with the distance duality relation, the relationship in $\mathcal{F}(z)$ can be re-expressed to include BAO measurements as
$\mathcal{F}(z)=\frac{D_{l}^{GW}(z)\theta_{BAO}(z)}{(1+z)r_{s}}.$ (2.7)
The ratio of the comoving sound horizon at the drag epoch to that at the recombination epoch, obtained from CMB observations, helps refine this equation. This multi-messenger approach not only tests the validity of GR and $\Lambda$CDM but also examines potential deviations from the predicted acceleration of the universe under different models.
@@
This method, including both non-parametric and parametric forms, facilitates a thorough investigation of deviations from GR and dark energy’s equation of state across various redshifts. Future sections will detail results for $\mathcal{F}(z)$ both independently and in conjunction with the Hubble constant, allowing for adjustments based on the sound horizon’s precise estimation. GWs provide a unique method for measuring luminosity distances in binary systems, denoted as $\mathrm{D_{l}^{GW}}$. While standard $\Lambda$CDM cosmology also estimates these distances, it depends on model-specific assumptions. In contrast, for a model-independent, data-driven inference of luminosity distance, we utilize calculations of EM luminosity distance derived from various length scales.
@@
Data-driven inference of EM luminosity distance: We employ a data-driven methodology to determine the EM luminosity distance, which leverages measurements of the angular diameter distance, $D^{\rm EM}A(z)$, obtained from the angular peak position of the BAO ($\theta_{\rm BAO}(z)$) observed in large scale structures across various redshifts [60]. This approach utilizes the distance duality relation, which links the angular diameter distance $D^{\rm EM}_{a}(z)$ with the EM luminosity distance $D^{\rm EM}_{l}(z)$ according to the relation $D^{\rm EM}_{a}(z)=D^{\rm EM}_{l}(z)/(1+z)^{2}$. The BAO peak position, as identified in the galaxy two-point correlation function, provides a direct observational basis for this calculation. The associated angular scales connect to the comoving sound horizon up to the redshift at the drag epoch ($z_{d}\sim 1020$) through the relationship
$r_{s}=\int_{z_{d}}^{\infty}\frac{dz\,c_{s}(z)}{H_{0}\sqrt{\Omega_{m}(1+z)^{3}+\Omega_{\Lambda}+\Omega_{r}(1+z)^{4}}}.$ (2.8)
@@
Substituting $D_{l}^{\rm EM}$ in Equation (2.4) with $\theta_{\rm BAO}(z)$ and applying the distance duality relation for EM observations, we define the function [61, 56]
@@
Here, the comoving sound horizon at the drag epoch, $r_{s}$, is approximated to be about 1.02 times the comoving sound horizon at the redshift of recombination ($z_{*}=1090$), which is $r_{d}\approx 1.02r_{*}$. This factor $r_{*}$ is derived from CMB observations, specifically from the position of the first peak in the CMB temperature power spectrum, expressed as $\theta_{*}=r_{*}/(1+z_{*})D^{EM}_{a}(z_{*})$ [62, 63, 64].
In this framework, $D_{l}^{GW}(z)$ is determined from GWs events, $\theta_{\rm BAO}$ is extracted from galaxy two-point correlation functions, and $r_{s}$ is deduced from CMB temperature fluctuations. The redshift of GWs sources can be inferred from their EM counterparts, typically observable in events involving bright standard sirens. Consequently, the combination of these measurable quantities should conform to a consistent value across all redshifts if GR and the standard $\Lambda$CDM cosmological model hold true. However, any deviations from these established theories could be identified through changes in these measurements across different redshifts. It is crucial to note that while $r_{s}$ is model-dependent and varies above $z_{d}=1020$, it remains constant at lower redshifts. Thus, any errors in the determination of $r_{s}$ would affect the absolute scale of $\mathcal{F}(z)$, but not its redshift-dependent trend.
Future discussions will present results considering $\mathcal{F}(z)$ both independently and jointly with the Hubble constant $H_{0}$, allowing for corrections based on potential inaccuracies in the estimated sound horizon due to erroneous Hubble constant values. This analysis supports the exploration of deviations from GR and the canonical dark energy equation of state ($w=-1$) using a variety of observational probes. Previous studies [61] have explored parametric deviations from GR; however, our non-parametric approach using $\mathcal{F}(z)$ enables a detailed, redshift-specific reconstruction of deviations from GR through multi-messenger observations.

### v2

# A Model-Independent Precision Test of General Relativity using LISA Bright Standard Sirens

@@ Abstract
The upcoming Laser Interferometer Space Antenna (LISA), set for launch in the mid-2030s, will enhance our capability to probe the universe through gravitational waves (GWs) emitted from binary black holes (BBHs) across a broad range of cosmological distances. LISA is projected to observe three classes of BBHs: massive BBHs (MBBHs), extreme mass-ratio inspirals (EMRIs), and stellar mass BBHs. This study focuses on MBBHs, which are anticipated to occur in gas-rich environments conducive to producing powerful electromagnetic (EM) counterparts, positioning them as excellent candidates for bright sirens. By combining GW luminosity distance measurements from these bright sirens with Baryon Acoustic Oscillation (BAO) measurements derived from galaxy clustering and sound horizon measurements from the Cosmic Microwave Background (CMB), and spectroscopic redshift measurements from observations of the electromagnetic (EM) counterpart, we propose a data-driven model-independent method to reconstruct deviations in the variation of the effective Planck mass (in conjunction with the Hubble constant) as a function of cosmic redshift. Using this multi-messenger technique, we achieve precise measurements of deviations in the effective Planck mass variation with redshift (z), with a precision ranging from approximately $2.4\%$ to $7.2\%$ from redshift $z=1$ to $z=6$ with a single event. Additionally, we achieved a measurement of the Hubble constant with a precision of about $1.3\%$, accounting for variations in the effective Planck mass over 4 years of observation time ($T_{\mathrm{obs}}$). This assumes that EM counterparts are detected for $75\%$ of the events. This precision improves with observation time as $T_{\mathrm{obs}}^{-1/2}$. This approach not only has the potential to reveal deviations from General Relativity but also to significantly expand our understanding of the universe’s fundamental physical properties.
@@ 1 Introduction
However, it is the forthcoming space-based observatory, the Laser Interferometer Space Antenna (LISA)[17, 18, 19], that is set to revolutionize our understanding by observing GWs from binary black holes (BBHs) over unprecedented cosmological distances. The mission has recently been adopted by ESA with a possible launch by 2036.
@@
LISA is designed to explore the milliHertz frequency range, which is rich with sources such as massive BBHs (MBBHs) [20], extreme mass ratio inspirals (EMRIs) [21, 22], and stellar mass BHBs [23] each with unique observational characteristics and expected in different redshift ranges. This ability to probe such diverse astrophysical phenomena is crucial because ground-based detectors are hindered by unavoidable noise artifact, which prevents them from accessing this lower frequency domain. Scheduled for launch in the mid-2030s, LISA will detect these sources up to very high redshifts with unparalleled precision [24]. Its ability to observe with unique observational characteristics across different redshift ranges positions it as a pivotal cosmological tool [25]. This extensive range allows for the mapping of the universe’s expansion from local to very distant realms. Notably, MBBHs, which are anticipated to form in gas-rich environments, may generate significant electromagnetic (EM) counterparts through various processes such as jets, disk winds, or accretion [26, 27, 28, 29, 30, 31]. The detection of these EM emissions not only enables precise localization of MBH events in the sky but also makes them exceptionally valuable as bright sirens for gravitational studies. The combination of GWs and EM observations could lead to breakthroughs in multi-messenger astronomy, further enhancing our understanding of the universe’s structure and evolution.
@@
Along with these, the investigation of modified gravity theories can gain insights through the propagation speed and luminosity distance of GWs and EM signals. Various factors, including the mass of gravitons, a frictional term related to the changing Planck mass, and anisotropic source terms, influence deviations from expected results [32, 33, 34, 35, 36, 37, 38]. Accurate measurements of these deviations provide a framework for rigorously testing alternate gravitational theories. Utilizing bright sirens is a particularly effective method for evaluating parameters like the GWs propagation speed, the mass of gravitons, and frictional effects denoted by $\mathrm{\gamma(z)}$ (see in Equation.2.2). The observation of an EM counterpart roughly 1.7 seconds following the binary neutron star merger GW170817 enabled the establishment of stringent limits on the speed of GWs and the graviton’s mass [6, 39, 40]. Some theories suggest variations in GWs speed, especially within frequency bands monitored by LISA, contrasting with those covered by LIGO [41]. For the purpose of this study, we assume that the speed of GWs is equivalent to that of light.
@@
The literature offers various model-dependent parameterizations of the friction term $\mathrm{\gamma(z)}$, including the popular ($\mathrm{\Xi_{0},n}$) model [36], which has been shown to be inferable with LVK [42, 43] and LISA [44]. This particular parameterization utilizes two positive parameters, $\mathrm{\Xi_{0}}$ and $\mathrm{n}$, where a $\mathrm{\Xi_{0}}$ value of 1 aligns with GR. Various models of modified gravity propose different behaviors for this frictional term under diverse conditions. Scalar-Tensor theories, for example, introduce a scalar field that significantly influences cosmological dynamics. Notable within this category are Brans–Dicke theory [45], f(R) gravity [46], and covariant Galileon models [47], all part of the broader Horndeski framework. Extending beyond Horndeski, Degenerate Higher Order Scalar-Tensor (DHOST) theories [48, 49, 50] represent a comprehensive category of scalar-tensor models that propagate a solitary scalar degree of freedom along with the massless graviton’s helicity-2 mode. Other notable theories include f(Q) gravity, f(T) gravity [51], minimal bigravity [52], and models incorporating extra dimensions [53, 54, 55].
@@
These theories demonstrate varying deviations from GR that are not fully captured by the power-law $\mathrm{\Xi_{0}}$ and $\mathrm{n}$ parameterization, particularly at the extremes of the redshift spectrum, where the fitting formula for $\mathrm{\gamma(z)}$ may lose accuracy [32]. Moreover, the integrated effect of $\mathrm{\gamma(z)}$ in the real universe may deviate from a simple power-law, underscoring the necessity for model-independent methods to test GR and measure $\mathrm{\gamma(z)}$ based on observational data. To address these issues, we propose a novel model-independent approach through a function designated as $\mathrm{\mathcal{F}(z)}$, defined in Equation 2.4. This function is crafted to quantify any departures from GR by comparing GWs luminosity distances $\mathrm{D_{L}^{GW}(z)}$ with their EM counterparts $\mathrm{D_{L}^{EM}(z)}$. In the context of standard GR, $\mathcal{F}(z)$ should theoretically equal 1, providing a clear metric to gauge discrepancies attributable to alternative theories of gravity.
@@
Figure 1: This diagram outlines the methodology for measuring three critical length scales: the sound horizon ($\mathrm{r_{s}}$), the BAO scale ($\mathrm{\theta_{BAO}}$), and the GW luminosity distance ($\mathrm{D_{L}^{GW}}$). These scales are derived from distinct observational methods, including CMB measurements, galaxy correlation studies, and detections of GWs. The interplay between the sound horizon and the BAO scale establishes the EM angular diameter distance ($\mathrm{D^{EM}_{a}}$). Utilizing the distance duality relation, these measurements facilitate a redshift-dependent empirical verification of GR, as expressed in Equation 2.7. This schematic provides a visual framework for understanding the multi-messenger approach in extracting these scales using LISA’s bright sirens.
@@ 2 Non-parametric reconstruction of deviation from General Relativity: Formalism
[…]
@@
where $\mathrm{\gamma(z)}$ represents a frictional term reflecting the influence of modified dynamics, $\mathrm{c_{GW}}$ denotes the speed of GWs propagation, $\mathrm{m_{GW}}$ is the graviton mass, $\mathrm{a}$ is the scale factor, and $\mathrm{\Pi_{(+,\times)}}$ is the anisotropic stress term.
Under standard GR assumptions, all the additional parameters such as $\mathrm{\gamma(z)}$, $\mathrm{m_{GW}}$, and $\mathrm{\Pi_{(+,\times)}}$ are set to zero, except for $\mathrm{c_{GW}}$, which is equal to the speed of light (denoted by $\mathrm{c}$). When testing the deviations from GR, these parameters are crucial to ascertain any potential discrepancies. Observations from the GW170817 event have placed strict limits, confirming that the graviton mass ($\mathrm{m_{GW}}$) is effectively zero, and that the speed of GWs propagation ($\mathrm{c_{GW}}$) aligns with the speed of light [39, 40]. Although some theories suggest variations in the speed of GWs outside the LVK observational frequency range [41], for the purpose of this analysis, we maintain the assumption that the speed of GWs is equivalent to that of light. Furthermore, changes in the anisotropic stress term ($\mathrm{\Pi_{+,\times}}$) can affect the phase of binary waveforms. Recent observations, especially of GW events like GW150914 and GW151226, have begun to constrain these potential deviations, although the constraints remain relatively broad [56]. The friction term, denoted as $\mathrm{\gamma(z)}$, significantly affects the amplitude of GWs from cosmological sources. This alteration suggests that the luminosity distance to GWs, measured via standard sirens, could differ from distances determined through EM probes like standard candles, CMB, or BAO. This difference potentially serves as a distinct indicator of modified gravity theories. Hence, the GWs luminosity distance, $\mathrm{D_{L}^{GW}}$, is linked to the EM luminosity distance, $\mathrm{D_{L}^{EM}}$, through an exponential relationship with the exponent being the integral of $\mathrm{\gamma(z^{\prime})/(1+z^{\prime})}$ [35]
@@
here, $\mathcal{F}(z)$ encapsulates any deviations from GR as a function of redshift, aiding in the identification of modified gravity effects through observed GWs and EM luminosity distances introduced in the paper[57] . Alternatively, a parametric approach that employs $\mathrm{\Xi_{0}}$ and $\mathrm{n}$ is used to describe a power-law modification in GW propagation over redshift[58, 59, 44]
@@
This method, including both non-parametric and parametric forms, facilitates a thorough investigation of deviations from GR and dark energy’s equation of state across various redshifts. Future sections will detail results for $\mathcal{F}(z)$ both independently and in conjunction with the Hubble constant, allowing for adjustments based on the sound horizon’s precise estimation. GWs provide a unique method for measuring luminosity distances in binary systems, denoted as $\mathrm{D_{L}^{GW}}$. While standard $\mathrm{\Lambda}$CDM cosmology also estimates these distances, it depends on model-specific assumptions. In contrast, for a model-independent, data-driven inference of luminosity distance, we utilize calculations of EM luminosity distance derived from various length scales.
Data-driven inference of EM luminosity distance: We employ a data-driven methodology to determine the EM luminosity distance, which leverages measurements of the angular diameter distance, $\mathrm{D^{EM}_{a}(z)}$, obtained from the angular peak position of the BAO ($\mathrm{\theta_{BAO}(z)}$) observed in large scale structures across various redshifts [60]. This approach utilizes the Etherington’s distance duality relation, which establishes a fundamental connection between the EM angular diameter distance $\mathrm{D^{EM}_{a}(z)}$ and the EM luminosity distance $\mathrm{D^{EM}_{L}(z)}$. The distance duality relation is given by: $\mathrm{D^{EM}_{L}(z)=(1+z)^{2}D^{EM}_{a}(z)}$. This relation is derived from the fact that in a universe described by the cosmological principle (isotropy and homogeneity), the amount of light emitted from a source spreads out over an increasing area as it travels through the expanding universe. The EM angular diameter distance $\mathrm{D^{EM}_{a}(z)}$ measures how large an object appears on the sky for a given physical size, while the EM luminosity distance $\mathrm{D^{EM}_{L}(z)}$ relates to how bright an object appears based on its intrinsic luminosity. The factor of $\mathrm{(1+z)^{2}}$ arises because of the cosmological redshift $\mathrm{z}$, which affects both the wavelength of the photons and the flux of the light received. Specifically, one factor of $\mathrm{(1+z)}$ accounts for the redshift stretching the wavelength, thereby reducing the energy per photon, and the second factor of $\mathrm{(1+z)}$ accounts for the time dilation effect, which reduces the rate at which photons are received. Thus, this relation allows us to convert measurements of the EMA angular diameter distance, which are typically easier to obtain from observations such as the BAO, into the EM luminosity distance needed for cosmological analyses.
@@
The BAO peak position, as identified in the galaxy two-point correlation function, provides a direct observational basis for this calculation. The associated angular scales connect to the comoving sound horizon up to the redshift at the drag epoch ($\mathrm{z_{d}\sim 1020}$) through the relationship
$\mathrm{r_{s}=\int_{z_{d}}^{\infty}\frac{dz\,c_{s}(z)}{H_{0}\sqrt{\Omega_{m}(1+z)^{3}+\Omega_{\Lambda}+\Omega_{r}(1+z)^{4}}}}.$ (2.6)
@@
Substituting $\rm{D_{L}^{EM}}$ in Equation (2.4) with $\theta_{\rm BAO}(z)$ and applying the distance duality relation for EM observations, we define the function [61, 57]
@@
Here, the comoving sound horizon at the drag epoch, $\mathrm{r_{s}}$, is approximated to be about 1.02 times the comoving sound horizon at the redshift of recombination ($\mathrm{z_{*}=1090}$), which is $\mathrm{r_{d}\approx 1.02r_{*}}$. This factor $\mathrm{r_{*}}$ is derived from CMB observations, specifically from the position of the first peak in the CMB temperature power spectrum, expressed as $\mathrm{\theta_{*}=r_{*}/(1+z_{*})D^{EM}_{a}(z_{*})}$ [62, 63, 64].
In this framework, $\mathrm{D_{L}^{GW}(z)}$ is determined from GWs events, $\mathrm{\theta_{BAO}}$ is extracted from galaxy two-point correlation functions, and $\mathrm{r_{s}}$ is deduced from CMB temperature fluctuations. The redshift of GWs sources can be inferred from their EM counterparts, typically observable in events involving bright standard sirens. This methodology is illustrated by a schematic diagram in Figure 1. Consequently, the combination of these measurable quantities should conform to a consistent value across all redshifts if GR and the standard $\mathrm{\Lambda}$CDM cosmological model hold true. However, any deviations from these established theories could be identified through changes in these measurements across different redshifts. It is crucial to note that while $\mathrm{r_{s}}$ is model-dependent and varies above $\mathrm{z_{d}=1020}$, it remains constant at lower redshifts. Thus, any errors in the determination of $\mathrm{r_{s}}$ would affect the absolute scale of $\mathrm{\mathcal{F}(z)}$, but not its redshift-dependent trend.
Future discussions will present results considering $\mathrm{\mathcal{F}(z)}$ both independently and jointly with the Hubble constant $\mathrm{H_{0}}$, allowing for corrections based on potential inaccuracies in the estimated sound horizon due to erroneous Hubble constant values. This analysis supports the exploration of deviations from GR and the canonical dark energy equation of state ($\mathrm{w=-1}$) using a variety of observational probes. Previous studies [61] have explored parametric deviations from GR; however, our non-parametric approach using $\mathrm{\mathcal{F}(z)}$ enables a detailed, redshift-specific reconstruction of deviations from GR through multi-messenger observations.